# An End-to-End Model for Photo-Sharing Multi-modal Dialogue Generation

### v1

@@ 2.4.2. Dynamic Vocabulary Transformation Matrix
Aiming to achieve the end-to-end photo-sharing multi-modal dialogue model, we can not convert the sampled one-hot output $\bm{R}^{\text{LLM}}=(\bm{r}_{1},\cdots,\bm{r}_{m})$ into textual caption by large language model tokenizer and vocabulary, otherwise the gradient propagation will be interrupted. Therefore we need to feed the vectorized output into the stable diffusion model directly, which will not lose differentiability. However, the vocabularies of the large language model $V^{\text{LLM}}$ and the stable diffusion model $V^{\text{SD}}$ are mismatched. To align the mismatched vocabularies, we propose vocabulary transformation matrix $\bm{M}$ to implement the conversion of one-hot character representations, which is defined as follows:
(8) $\bm{M}_{i,j}=\begin{cases}1,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{if}~{}~{}V^{\text{LLM}}[i]=V^{\text{SD}}[j],\\ 0,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{otherwise,}\end{cases}$
where $V^{\text{LLM}}[i]$ and $V^{\text{SD}}[j]$ indicate the $i$-th and $j$-th token in the vocabulary of large language model and stable diffusion model, respectively. Based on the vocabulary transformation matrix $\bm{M}$, we can acquire one-hot image text caption input with the format of one-hot vectorized representation $\bm{R}^{\text{SD}}$ for the image generation model:
(9) $\bm{R}^{\text{SD}}$ $=(\bm{r}_{1}^{\text{SD}},\cdots,\bm{r}_{n}^{\text{SD}})$
$=(\bm{r}_{1}^{\text{LLM}}\bm{M},\cdots,\bm{r}_{n}^{\text{LLM}}\bm{M})$
$=\bm{R}^{\text{LLM}}\bm{M}$
In other words, we formalize the mapping of one-hot image text caption representation from $V^{\text{LLM}}$ to $V^{\text{SD}}$ as matrix multiplication. Besides, the operation can also ensure gradient propagation between different vocabularies.
@@
(10) $\bm{M}_{i,j}=\begin{cases}1,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{if}~{}~{}V^{\text{LLM}}[i]\in\mathcal{T}^{\text{LLM}}\bigwedge V^{\text{SD}}[j]\in\mathcal{T}^{\text{SD}},\\ 0,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{otherwise,}\end{cases}$
@@
During the training phase, we initialize the vocabulary transformation matrix for each image caption in the batch dialogue. It is important to note that the vocabulary transformation matrix can lead to significant GPU memory overhead and limited batch size. Suppose the sizes of the large language model and stable diffusion model vocabularies are both 40,000, each matrix will occupy about $40,000*40,000*2=1,600,000,000*2=3.2G$ bytes GPU space. However, the transformation matrix is a sparse matrix factually. We hence implement it through the special sparse tensor class in numpy (Harris et al., 2020) and pytorch (Ansel et al., 2024) [^1] https://pytorch.org/docs/stable/sparse.html, which constrains GPU memory occupation to a magnitude of kilobytes (KB).
@@
We acquire vector representation of the image text caption on the stable diffusion model vocabulary according to the proposed sentence-level dynamic vocabulary transformation matrix and Equation 9. However, the vector representation is not a standard one-hot vector format. All positions of image caption tokens in each vector are 1, not only one position. In addition, another intractable problem is that the length of the converted image caption is not adaptive to the tokenization of the stable diffusion model. To address these issues, we first use the textual image caption to get a valid one-hot vector representation $\widetilde{\bm{R}^{\text{SD}}}$ with the correct length by the stable diffusion model tokenizer. Second, we feed the converted representation $\bm{R}^{\text{SD}}$ into average pooling to eliminate sequence length dimension. Third, straight-through gradient estimator (Bengio et al., 2013) is adopted to propagate gradients again:
(11) $\bm{R}^{\text{SD}}$ $=\widetilde{\bm{R}^{\text{SD}}}-\text{sg}[\text{avg}(\bm{R}^{\text{LLM}}\bm{M})]+\text{avg}(\bm{R}^{\text{LLM}}\bm{M})$
@@
where avg indicates average pooling. For inference, we directly feed $\widetilde{\bm{R}^{\text{SD}}}$ to the stable diffusion model because gradients is not necessary.
2.4.3. Stable Diffusion Model
The stable diffusion model (Rombach et al., 2022) is employed in our end-to-end photo-sharing multi-modal dialogue system for image generation. Concretely, the stable diffusion model consists of a clip text encoder (Radford et al., 2021) and an image U-Net model (Ronneberger et al., 2015). The former encodes the representation of image text captions, which is feed to the U-Net model. The latter iteratively performs the denoising process on the sampled gaussian noise $\bm{a}$ to generate images, guided by the text representations of image text descriptions from the clip text encoder. Finally, we exploit mean squared error (MSE) loss between the predicted image $\widehat{\bm{I}}$ and the gold standard image $\bm{I}$ to train the model:
(12) $\widehat{\bm{I}}$ $=\text{StableDiffusion}(\bm{R}^{\text{SD}},\bm{a})$
$\mathcal{L}^{\text{v}}$ $=\|\widehat{\bm{I}}-\bm{I}\|$
@@ 2.4.3. Stable Diffusion Model
2.5. End-to-End Training
For the input end, Q-Former connects text and vision modality, which gives the large language model the ability to see. For the output end, the dynamic vocabulary transformation matrix integrates Straight-Through Gumbel-Softmax to implement end-to-end gradient propagation, enabling the large language model to draw images. Our end-to-end photo-sharing multi-modal dialogue realizes fluent gradient flow throughout the entire dialogue system, which can be jointly trained:
(13) $\mathcal{L}$ $=\mathcal{L}^{\text{t}}+\alpha\mathcal{L}^{\text{v}}$
where $\alpha$ is the weighted coefficient.
For the inference stage, the image perception encodes all the photos and then feeds visual representations to the text dialogue generator as dialogue contexts. The large language model generates the text response, where contents between image start token [IMG] and image end token [/IMG] are directly inputted to the stable diffusion model for quick image generation.
3. Experiments
We conduct experiments to verify the effectiveness of our proposed end-to-end model for photo-sharing multi-modal dialogue. In this section, we start with an introduction of experimental settings, then show main experiment results to verify the effectiveness of our end-to-end multi-modal dialogue model and end with some fine-grained experimental analyses to gain a deeper understanding of our approach.
3.1. Experimental Setup
3.1.1. Datasets
We conduct experiments on two datasets: PhotoChat (Zang et al., 2021) [^2] https://github.com/google-research/google-research/tree/master/multimodalchat/ and DialogCC (Lee et al., 2024) [^3] https://github.com/passing2961/DialogCC. We obtain text dialogue contents from corresponding GitHub links. For shared photos in the datasets, we download them according to URLs of images. However, some image are not accessible now due to expired URLs. We leverage stable diffusion 3 (Esser et al., 2024) [^4] https://huggingface.co/stabilityai/stable-diffusion-3-medium-diffusers to generate the missing images based on the image text captions in datasets. We employ standard train, dev and test dataset splits for model training, evaluation and testing, respectively. Table 1 lists dataset statistics. We report nine statistics from left to right: “the name of datasets”, “splits of datasets”, “the number of dialogues”, “the number of utterances”, “the number of utterances when counting consecutive utterances from the same speaker as one utterance”, “average consecutive utterances per dialogue”, “the number of images”, “the number of distinct images” and “average images per dialogue”. In particular, utterances in which photos are shared are excluded in the calculation.
Table 1. Dataset statistics. From left to right beginning from the third column, the name of each column means “the number of dialogues”, “the number of utterances”, “the number of utterances when counting consecutive utterances from the same speaker as one utterance”, “average consecutive utterances per dialogue”, “the number of images”, “the number of distinct images” and “average images per dialogue”. Utterances in which photos are shared are excluded in the calculation.
Dataset Split # Dial # Utter # CUtter U/D # Img # DImg I/D
PhotoChat (Zang et al., 2021) train 10,286 130,459 92,479 8.99 10,286 8,889 1.0
dev 1,000 12,695 9,045 9.05 1,000 1,000 1.0
test 1,000 12,841 9,089 9.09 1,000 1,000 1.0
total 12,286 155,995 110,613 9.00 12,286 10,889 1.0
DialogCC (Lee et al., 2024) train 68,402 553,990 329,969 4.82 106,198 38,743 1.55
dev 7,644 63,140 37,181 4.86 11,669 6,413 1.53
test 7,324 60,260 35,633 4.87 11,157 6,279 1.52
total 83,370 677,390 402,783 4.83 129,024 51,435 1.55
3.1.2. Hyperparameters
For the text dialogue generator, we exploit Llama-2-7b-chat-hf (Touvron et al., 2023) as the dialogue backbone. For the image perceptron, the ViT encoder and Q-Former are directly derived from the pre-trained BLIP-2 (blip2-opt-2.7b-coco) (Li et al., 2023). For the image generator, we fine-tune the stable-diffusion-2-1-base (Rombach et al., 2022), which consists of the pre-trained CLIP encoder (Radford et al., 2021) and U-Net model (Ronneberger et al., 2015). Besides, for hyper-parameters of model training, we use the AdamW algorithm with learning rate 5e-5, batch size 32, weight decay 0.01 and linear learning rate warmup over the first 1,000 steps to optimize parameters. The temperature parameter $\tau$ is gradually annealed from 1 to 0.01 in the first three epochs and remains constant in the subsequent rounds. Weight factor $\alpha$ is 1. We stop early training when the BLEU-2 and FID scores do not increase on the development set for 4 epochs. We run all experiments on Nvidia V100.
3.1.3. Evaluation
We evaluate our end-to-end model on both text and image generation for the photo-sharing multi-modal dialogue task. For text generation, we leverage BLEU-1, BLEU-2 (Papineni et al., 2002) and Rouge-L (Lin, 2004) as evaluation metrics. For image generation, we follow Sun et al. (2022) to use Inception Score (IS) (Salimans et al., 2016) and Frechet Inception Distance (FID) (Heusel et al., 2017), which are computed by torch-fidelity  [^5] https://github.com/toshas/torch-fidelity toolkit. We conduct the experiments on three different random seeds and report the average results.
3.1.4. Baselines
To verify the effectiveness of our end-to-end dialogue model, we compare results with two related baseline models:
  • Divter: Sun et al. (2022) use DialoGPT (Zhang et al., 2020) and DALL-E (Ramesh et al., 2021) to implement the pipeline-based model, where discrete image text captions bridge the text and image generation models.
  • pipeline: We implement the pipeline model for a fair comparison with our end-to-end model. This model adapts the same Llama-2 and stable-diffusion-2-1 models as the end-to-end model, but trains them separately.
3.2. Main Results
Table 2. Main results for text and image response generation. indicates the model is trained on extra resources of image generation.
Model Text Response Generation Image Response Generation
BLEU-1 BLEU-2 Rouge-L IS $\uparrow$ FID $\downarrow$
PhotoChat
Divter (Sun et al., 2022) 6.52 1.66 5.69 15.8 $\pm$ 0.6 29.16
Divter (Sun et al., 2022) 6.28 1.51 5.40  4.9 $\pm$ 0.7 262.09
pipeline 10.65 2.03 9.76  13.67 $\pm$ 1.31 79.74
end-to-end 12.08 2.94 11.00  14.47 $\pm$ 0.94 75.88
DialogCC
Divter (Sun et al., 2022) 8.08 1.89 6.71  9.19 $\pm$ 2.01 108.50
pipeline 14.18 2.93 12.21  22.99 $\pm$ 1.92 61.66
end-to-end 15.84 3.97 13.70 25.56 $\pm$ 1.79 59.63
Table 2 reports the main results of different models, including BLEU-1, BLEU-2 and Rouge-L for text response generation and IS and FID for image response generation.
First, we examine the performances of the pipeline and end-to-end model. For the PhotoChat dataset, our end-to-end model can bring significant improvements for photo-sharing multi-modal dialogue, where gains are $12.08-10.65=1.43$, $2.94-2.03=0.91$ and $11.00-9.76=1.24$ on BLUE-1, BLUE-2 and Rouge-L for text response generation, and $14.47-13.67=0.80$ and $79.74-75.88=3.86$ on IS and FID for image response generation. For the DialogCC dataset, the end-to-end model also advances the text response generation scores by $15.84-14.18=1.66$, $3.97-2.93=1.04$ and $13.70-12.21=1.49$ on BLUE-1, BLUE-2 and Rouge-L, and the image response generation scores by $25.56-22.99=2.57$ and $61.66-59.63=2.03$ on IS and FID, respectively. All enhancements verify the effectiveness of our proposed end-to-end photo-sharing multi-modal dialogue model compared with the basic pipeline model.
Second, we make comparisons between our end-to-end model and Divter, which is also a representative pipeline model. For the DialogCC dataset, our proposed end-to-end model significantly outperforms Divter, which might be due to the strong capabilities of text generation by Llama-2-7b-chat-hf and image generation by stable-diffusion-2-1-base. Concretely, the end-to-end model boosts the performances of text and image response generation significantly, leading to increases on BLUE-1 by $15.84-8.08=7.76$, BLUE-2 by $3.97-1.89=2.08$, Rouge-L by $13.70-6.71=6.99$, IS by $25.56-9.19=16.37$ and FID by $108.50-59.63=48.87$, respectively. For the PhotoChat dataset, our end-to-end model still achieves better performances compared with naive Divter, leading improvements of text response generation on BLEU-1 by $12.08-6.28=5.80$, BLEU-2 by $2.94-1.51=1.43$ and Rouge-L by $11.00-5.40=5.60$, and image response Generation on IS by $14.47-4.9=9.57$ and FID by $262.09-75.88=186.21$, respectively. Our end-to-end model reaches state-of-the-art performance compared with pipeline models, which verifies the effectiveness of our proposed dialogue system.
Third, we look at the image response generation scores. The best results of IS and FID are $15.8$ and $29.16$, which is from Divter. An important reason could be that Sun et al. (2022) employ extra image caption resources to train the model, including YFCC100M (Thomee et al., 2016) and ImageNet (Deng et al., 2009). Sampling data from the two datasets might significantly influence the scores of IS and FID, because they are computed by the neural network trained on datasets with similar data distribution.
4. Conclusion

### v2

@@ III-D2 Dynamic Vocabulary Transformation Matrix
Aiming to achieve the end-to-end model, we can not convert the sampled one-hot output $\bm{R}^{\text{LLM}}$ into textual caption by LLM tokenizer and vocabulary, otherwise the gradient propagation will be interrupted. Therefore we need to feed the vectorized output into the stable diffusion model directly. However, the vocabularies of the LLM $V^{\text{LLM}}$ and the stable diffusion model $V^{\text{SD}}$ are mismatched. To align the mismatched vocabularies, we propose vocabulary transformation matrix $\bm{M}$ to implement the conversion of one-hot character representations, which is defined as follows:
$\bm{M}_{i,j}=\begin{cases}1,~{}~{}~{}~{}~{}~{}~{}\text{if}~{}~{}V^{\text{LLM}}[i]=V^{\text{SD}}[j],\\ 0,~{}~{}~{}~{}~{}~{}~{}\text{otherwise,}\end{cases}$ (3)
where $V^{\text{LLM}}[i]$ and $V^{\text{SD}}[j]$ indicate the $i$-th and $j$-th token in the vocabulary of LLM and stable diffusion model, respectively. Based on $\bm{M}$, we formalize the mapping of one-hot image text caption representation from $V^{\text{LLM}}$ to $V^{\text{SD}}$ as matrix multiplication (i.e., $\bm{R}^{\text{SD}}=\bm{R}^{\text{LLM}}\bm{M}$), which can ensure gradient propagation between different vocabularies.
@@
$\bm{M}_{i,j}=\begin{cases}1,~{}~{}\text{if}~{}~{}V^{\text{LLM}}[i]\in\mathcal{T}^{\text{LLM}}\bigwedge V^{\text{SD}}[j]\in\mathcal{T}^{\text{SD}},\\ 0,~{}~{}\text{otherwise,}\end{cases}$ (4)
@@
During the training phase, we initialize the vocabulary transformation matrix for each image caption in the batch dialogue. It is important to note that the vocabulary transformation matrix can lead to significant GPU memory overhead. Suppose the sizes of the LLM and stable diffusion model vocabularies are both 40,000, each matrix will occupy about $40,000*40,000*2=1,600,000,000*2=3.2G$ bytes GPU space. However, the transformation matrix is sparse factually. We hence implement it through the special sparse tensor class in numpy and pytorch, which constrains GPU memory occupation to a magnitude of kilobytes (KB).
@@
Although we acquire the vector representation of the image caption, it is not a standard one-hot vector format. All positions of image caption tokens in each vector are 1, not only one position. In addition, another intractable problem is that the length of the converted image caption is not adaptive to the tokenization of the stable diffusion model. To address these issues, we first use the textual image caption to get a valid one-hot vector representation $\widetilde{\bm{R}^{\text{SD}}}$ with the correct length by the stable diffusion model tokenizer. Second, we feed the converted representation $\bm{R}^{\text{SD}}$ into average pooling to eliminate sequence length dimension. Third, straight-through gradient estimator [13] is adopted to propagate gradients again:
$\bm{R}^{\text{SD}}$ $=\widetilde{\bm{R}^{\text{SD}}}-\text{sg}[\text{avg}(\bm{R}^{\text{LLM}}\bm{M})]+\text{avg}(\bm{R}^{\text{LLM}}\bm{M})$ (5)
@@
where avg indicates average pooling. Finally, we employ the stable diffusion model [12] for image generation based on $\bm{R}^{\text{SD}}$. For inference, we directly feed $\widetilde{\bm{R}^{\text{SD}}}$ to the stable diffusion model because gradients is not necessary.
III-E End-to-End Training
For the input end, MLLM connects text and vision modality, which gives the LLM the ability to see. For the output end, the dynamic vocabulary transformation matrix integrates Straight-Through Gumbel-Softmax to implement end-to-end gradient propagation, enabling the LLM to draw images. Our end-to-end model realizes fluent gradient flow throughout the entire dialogue system, which can be jointly trained:
$\mathcal{L}$ $=\mathcal{L}^{\text{t}}+\alpha\mathcal{L}^{\text{v}},$ (6)
where $\alpha$ is the weighted coefficient, and $\mathcal{L}^{\text{v}}$ is the loss of the stable diffusion model for image generation.
IV Experiments
IV-A Experimental Setup
IV-A1 Datasets
We conduct experiments on two datasets: PhotoChat [7] and DialogCC [10]. We obtain text dialogue contents from corresponding GitHub repos. For shared photos in the datasets, we download them according to the URLs of images. However, some image are not accessible now due to expired URLs. We leverage stable diffusion 3 [12] to generate the missing images based on the image text captions in datasets. We employ standard train, dev and test dataset splits for model training, evaluation and testing, respectively.
IV-A2 Hyperparameters
For the text dialogue generator and image perceptron in MLLM, we exploit Llama-Vision [4] as the backbone. For the image generator, we fine-tune the stable-diffusion-3-medium [12]. Besides, for hyper-parameters of model training, we use the AdamW algorithm with learning rate 5e-5, batch size 32, weight decay 0.01 and linear learning rate warmup over the first 1,000 steps to optimize parameters. The temperature parameter $\tau$ is gradually annealed from 1 to 0.0001 in the first three epochs and remains constant in the subsequent rounds. Weight factor $\alpha$ is 1. We run all experiments on Nvidia V100.
IV-A3 Evaluation
We evaluate our end-to-end model on both text and image generation for the photo-sharing multi-modal dialogue task. For text generation, we leverage BLEU-1, BLEU-2 and Rouge-L as evaluation metrics. For image generation, we use Inception Score (IS) and Frechet Inception Distance (FID). We conduct the experiments on three different random seeds and report the average results.
IV-A4 Baselines
To verify the effectiveness of our end-to-end model, we compare results with two related baseline models: 1) Divter [8] uses DialoGPT and DALL-E to implement the pipeline-based model. 2) pipeline adapts the same Llama-3 and stable-diffusion-3-medium as the end-to-end model, but trains them separately. 3) GPT-4 [25] and 4) Chameleon [23] are two representative end-to-end MLLMs.
IV-B Main Results
TABLE I: Main results for text and image response generation. indicates the model is trained on extra resources of image generation.
Model Text Generation Image Generation
BL-1 BL-2 Ro-L IS $\uparrow$ FID $\downarrow$
PhotoChat
Divter [8] 6.52 1.66 5.69 15.8 $\pm$ 0.6 29.16
Divter [8] 6.28 1.51 5.40 4.9 $\pm$ 0.7 262.09
GPT-4 [25] 5.78 1.02 4.97 12.66 $\pm$ 1.05 131.40
Chameleon [23] 8.62 1.73 7.88 6.04 $\pm$ 1.63 177.79
pipeline 10.68 2.10 9.74 13.71 $\pm$ 1.35 79.94
end-to-end 12.14 2.98 11.12 14.59 $\pm$ 0.91 75.84
DialogCC
Divter [8] 8.08 1.89 6.71 9.19 $\pm$ 2.01 108.50
GPT-4 [25] 6.13 0.87 4.04 11.50 $\pm$ 1.97 81.56
Chameleon [23] 11.84 2.07 8.32 10.08 $\pm$ 1.34 101.72
pipeline 14.21 2.95 12.32 23.11 $\pm$ 1.85 61.38
end-to-end 15.87 4.02 13.80 25.61 $\pm$ 1.66 59.40
TABLE I reports the main results of different models. First, we examine the performances of the pipeline and end-to-end model. For the PhotoChat dataset, our end-to-end model can bring significant improvements for photo-sharing multi-modal dialogue, where gains are $12.14-10.68=1.46$, $2.98-2.10=0.88$ and $11.12-9.74=1.38$ on BLUE-1, BLUE-2 and Rouge-L for text response generation, and $14.59-13.71=0.88$ and $79.94-75.84=4.10$ on IS and FID for image response generation. For the DialogCC dataset, the end-to-end model also advances the text response generation scores by $1.66$, $1.07$ and $1.48$ on BLUE-1, BLUE-2 and Rouge-L, and the image response generation scores by $2.50$ and $1.98$ on IS and FID, respectively. All enhancements verify the effectiveness of our proposed end-to-end model compared with the basic pipeline model.
Second, we make comparisons between our end-to-end model and Divter, which is also a representative pipeline model. For the DialogCC dataset, our end-to-end model significantly outperforms Divter, which might be due to the strong capabilities of text generation by Llama-3 and image generation by stable-diffusion-3. Concretely, the end-to-end model boosts the performances of text and image response generation significantly, leading to increases on BLUE-1 by $7.79$, BLUE-2 by $2.13$, Rouge-L by $7.09$, IS by $16.42$ and FID by $49.10$, respectively. For the PhotoChat dataset, our end-to-end model still achieves better performances compared with naive Divter. However, the best results of IS and FID are $15.8$ and $29.16$, which is from Divter. An important reason could be that [8] employs extra image caption resources to train the model. Sampling data from the two datasets might significantly influence the scores of IS and FID, because they are computed by the neural network trained on datasets with similar data distribution.
Third, we observe the results of other end-to-end multi-modal models. The performance of GPT-4 is limited, because it is a close-source model and could not be fine-tuned on specific datasets. We find its text responses are usually too long, but the ground-truths are oral expressions and short. For Chameleon, image generation results underperform our end-to-end model. The reason could be that Chameleon opens only the text transformer parameter weights, but masks image generation parameter weights.
IV-C Analyses
IV-C1 Ablation Experiment.
TABLE II: Ablation experiment.
Model Text Generation Image Generation
BL-1 BL-2 Ro-L IS $\uparrow$ FID $\downarrow$
end-to-end 15.87 4.02 13.80 25.61 $\pm$ 1.66 59.40
-Image Perceptron -1.36 -0.78 -1.18 -1.29 +0.70
-Image Generator -0.37 -0.21 -0.42 -1.94 +1.47
The core of our end-to-end model is to replace the discrete image text descriptions in the pipeline model with dense vector representations, therefore gradient can propagate throughout the whole model without interruption. We conduct an ablation experiment to show the influence of the proposed end-to-end techniques. TABLE II reports ablation experiment results. -Image Perceptron means the end-to-end model ablates dense visual representations from the image perceptron and feeds discrete image text captions into the text dialogue generator. -Image Generator means the text dialogue generator outputs discrete image text description for the diffusion model, which ablates the dynamic vocabulary transformation matrix and straight-through gumbel-softmax. The two ablations significantly decrease the text and image response generation results, verifying the effectiveness of our proposed end-to-end model.
IV-C2 Response Generation Performances for Different Speakers.
ppe2e13.2713.2713.2713.2714.7314.7314.7314.7315.415.415.415.417.0117.0117.0117.01BLUE-1ppe2e2.862.862.862.863.813.813.813.813.143.143.143.144.214.214.214.21BLUE-2ppe2e11.8911.8911.8911.8913.0713.0713.0713.0712.712.712.712.715.1215.1215.1215.12Rouge-Lpp speaker Ae2e speaker App speaker Be2e speaker Bppe2e21.9521.9521.9521.9524.2324.2324.2324.2323.6623.6623.6623.6628.1928.1928.1928.19ISppe2e65.1865.1865.1865.1863.6663.6663.6663.6657.3957.3957.3957.3955.0255.0255.0255.02FID
Figure 3: Response generation performances for different speakers.
There are two speakers in our photo-sharing multi-modal dialogue. We examine the text and image response generation performances of the pipeline (pp) and end-to-end (e2e) model for the two speakers respectively. Figure 3 illustrates the experimental results. Regardless of the pipeline model and end-to-end model, response performances of speaker B are superior to speaker A. The reason might be from the instruction tuning stage of large language models, where the loss function is computed on the user side. In our experiments, we employ the chat mode of large language models, and take the assistant side as speaker A, while the user side as speaker B. Additionally, Figure 3 also demonstrates that the end-to-end model advances all results compared with the pipeline model, verifying the effectiveness of our end-to-end model.
IV-C3 Image Response Generation Performance for Different Gumbel-Softmax Temperature.
@@ IV-C Analyses
Figure 4: Image response generation performance for different temperatures of gumbel-softmax.
The temperature of gumbel-softmax plays an important role in our end-to-end photo-sharing multi-modal dialogue generation model. We conduct an analysis experiment to evaluate the relation between the temperature and image generation performance. The results are demonstrated in Figure 4, where the x-axis denotes the temperature of gumbel-softmax and the y-axis denotes IS and FIS on the left and right, respectively. When the temperature is 1, gumbel-softmax is equivalent to standard softmax. As the temperature reduces lower from 1, the IS scores increase significantly, which verifies the effectiveness of gumbel-softmax for our end-to-end model. The performance stops increasing after $e^{-4}$, which could be due to the weakened randomness. There is a similar tendency for FID, but it is important to note that lower values are better for FID.
V Conclusion